Translate the following into a quadratic equation, and solve it:
The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are its dimensions?
I don't see how I'm going to get the x2 part of the equation out of this..or any of it. please help!

Answer by scott8148(6628)

let x=width so 3x=length ... area=width*length ... 75=x*3x ... 75=3x^2 ... 3x^2-75=0

dividing by 3 gives x^2-25=0 ... difference of two squares factors to (x+5)(x-5)=0

x=5 and x=-5 ... negative is not realistic ... dimensions are 5 and 15